Your firm is considering investing in a small stand-alone plant. It will be located on a site that the firm currently owns. The firm paid $2,000,000 for this site seven years ago. They recently spent $100,000 to upgrade the site, in preparation for the project under consideration. Recently, they had an offer from the county government to buy the location for $925,000. The building and equipment needed for the operation will cost $2,400,000. These assets will be depreciated over six years by the straight-line method down to a value of zero. The operational life of this plant is equal to the six-year depreciable life. After six years you estimate that you can sell the equipment for $750,000. Assume that any capital gains or losses associated with the termination of the project will be treated at the firm’s marginal tax rate. Your marketing department estimates that annual sales will be $2,150,000. The accounting and engineering departments estimate that annual fixed costs to support this level of sales will be $275,000, and that variable costs will be 35% of sales. You will need to maintain a level of $130,000 in net working capital to support this operation. The firm’s marginal income tax rate is 23%, and the required return on such investments is 11%.

(A) Estimate the appropriate after-tax value of the site where the operation is located.

(B) Is the cost for recent site renovation relevant?

(C) Estimate the annual operating cash flows this project will generate over its six-year life.
Answer #1

Answer (A):

Current value of site = Offer from the county government to buy the location for amount = $925,000

Adjusted basis of site = 2000000 + 100000 = $2,100,000

If sold tax benefit on loss = (925000 - 2100000) * 23% = $270,250

Appropriate after-tax value of the site where the operation is located = 925000 + 270250 = $1195250

Appropriate after-tax value of the site where the operation is located = $1,195,250

Answer (B):

As calculated it is not relevant since it is already incurred. However, this increases adjusted basis and as such relevant only for calculation of tax benefit.

Answer (C):

Annual depreciation = (Cost - Salvage value) / useful life = 2400000 / 6 = $400,000

Operating cash flow = (Sales - Variable cost - Cash Fixed costs) * (1 - Tax rate) + depreciation tax shield

= (2150000 - 2150000 * 35% - 275000) * (1 - 23%) + 400000 * 23%

= $956,325

Annual Operating cash flow = $956,325